OpenStudy (anonymous):

solve the equation in the real number system: x^4-8x^3+16x^2 +8x-17=0

OpenStudy (anonymous):

by inspection 1 is a solution

OpenStudy (anonymous):

because \[1-8+16+8-17=0\]

OpenStudy (anonymous):

so you know this factors as \[(x-1)\times \text{something}\] and you find the something by division

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

always check to see if 1 works because it is easy. -1 is easy too. check it

OpenStudy (anonymous):

then the answer is x^2-8x+17=0 right??

OpenStudy (anonymous):

so you know this factors as \[(x-1)(x^3-4x^2+9x+17)\] by division.

OpenStudy (anonymous):

you can use synthetic division to divide by x - 1

OpenStudy (anonymous):

sorry i made mistake.

OpenStudy (anonymous):

it is \[(x-1)(x^3-7x^2+9x+17)\]

OpenStudy (anonymous):

that is the correct answer. now check the second factor and you will see that -1 is a zero

OpenStudy (anonymous):

:).. ok...

OpenStudy (anonymous):

i am sorry about that and you are right! it is \[(x-1)(x+1)(x^2-8x+17)\] you had it all along!

OpenStudy (anonymous):

my mistake i should have read your answer more carefully. you got it! third factor has no real zeros so you are done

OpenStudy (anonymous):

you were right 6 lines up. done!

OpenStudy (anonymous):

x^2-8x+17=0 right?

OpenStudy (anonymous):

yes that is the last factor.

OpenStudy (anonymous):

this has no real zeros

OpenStudy (anonymous):

and since your instructions were "solve in real number system" you are done. real zeros are 1 and -1

OpenStudy (anonymous):

good work!